Nevey:

Make tables to find the solution to 2−x = 4x + 3. Take the integer values of x only between −3 and 3

Razor:

1 attachment
Nevey:

this isn't explaining this problem razor. Thank u, but this does not help me..

Razor:

._. Waittttttttttttt

Nevey:

?

Razor:

Is this the full question?

Nevey:

no its not. heres the full thing: Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2−x and y = 4x + 3 intersect are the solutions of the equation 2−x = 4x + 3. (4 points) Part B: Make tables to find the solution to 2−x = 4x + 3. Take the integer values of x only between −3 and 3. (4 points) Part C: How can you solve the equation 2−x = 4x + 3 graphically? (2 points)

Nevey:

1 attachment
Razor:

A. We have two lines: y = 2-x and y = 4x+3 Given two simultaneous equations that are both required to be true.. the solution is the points where the lines cross... Which is where the two equations are equal.. Thus the solution that works for both equations is when 2-x = 4x+3 because where that is true is where the two lines will cross and that is the common point that satisfies both equations.

Razor:

Thats part A.

Nevey:

alright. reading it.

Shadow:

Make two tables, one for each expression. Input values between -3 and 3 (labeled as x since they are inputs) and write down the result under y (as the result would be the output). Remember, we are just looking for where these graphs intersect (have the same output for the same input).

Nevey:

so razor, the simplified version of this could very well be 'Because at that x point, they give the same output (y value).', correct?

Razor:

Yep ._. and Shadow helped tew

Nevey:

alrighty thn lol. reading shadows now.

Nevey:

shadow, could i possible put them in desmos to do that or no?

Shadow:

I know that you can make tables in desmos, although I do not know how. Why do you wish to do so anyway?

Nevey:

because it says 'make tables' .-.

Shadow:

Is this online?

Nevey:

the question?

Shadow:

Online school

Nevey:

oh, yea. it is. y?

Shadow:

They don’t have a tool or something?

Nevey:

they have desmos on their homepage, which is didnt even notice until recently

Shadow:

Just do this: Table 1 x: 1 y: 3 etc

Shadow:

I mean if you can upload files, you can try create them there, and then see if there is a download feature, or just use a screenshot tool.

Nevey:

yea. i will try the screenshot thing first lol.

Shadow:

Okay well do you know what to do regarding inputting x values and getting y values

Nevey:

apparently not bc i just put in the equation and this is what i got. nada.

1 attachment
Shadow:

1. You did not input the equation in correctly. There are no exponents. 2. That is not how you solve this problem...

Shadow:

2 - x 2 - (-2) 2+ 2 = 4 Table 1 x = -2, y = 4

Shadow:

Then input -2 into the other expression and start table two. Then go all the way to positive 2...

Nevey:

so i reput it in, taking away the exponents and putting them in normally and this is what i got:

1 attachment
Shadow:

That is not how you solve this. Maybe try read what I posted.

Nevey:

i read it but i still dont get it.. :/

Shadow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow 2 - x 2 - (-2) 2+ 2 = 4 Table 1 x = -2, y = 4 \(\color{#0cbb34}{\text{End of Quote}}\)

Shadow:

Do you get what I did here?

Shadow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nevey Make tables to find the solution to 2−x = 4x + 3. Take the integer values of x only between −3 and 3 \(\color{#0cbb34}{\text{End of Quote}}\)

Shadow:

I am doing that in what I quoted.

Nevey:

no, i dont get what u did that. where did the -2 come from and crap, i just dont get it...

Nevey:

im sorry im terrible at this kind of thing...

Shadow:

An integer is a whole number. I took a whole number between -3 and 3 and inputted it into the first expression to see what I would get.

Nevey:

oh. ok. im starting to get it now.

Shadow:

Input -2 for the other expression, and start Table 2. Then keep inputting whole numbers til you get to 2

Shadow:

so -2, -1, 0, 1, and 2, are all your inputs

Nevey:

would it be 2 then?

Shadow:

what

Nevey:

i think im just gonna put in what i think best because this is just confusing me more and more. thank u for trying to help me tho. :)

Shadow:

Okay.